Project

General

Profile

Issues

Filters

Apply Clear

# Project Tracker Status Priority Subject Assignee Updated
954 SBS-DAQ-GEp Task New Normal ECAL testing during GEn RP 02/26/2024 11:17 AM Actions
950 NPS HMS Task New High HMS yield check by Peter 02/22/2024 09:06 AM Actions
949 SBS-DAQ-GEn-RP Bug In Progress High Ship 1190 for active analyzer David Hamilton 02/26/2024 11:03 AM Actions
948 SBS DAQ GEp CDet Task New Normal Make VETROC adapters 02/19/2024 08:06 AM Actions
947 SBS-DAQ-GEp Task New Normal GEM calorimeter based selective readout 02/19/2024 07:59 AM Actions
946 NPS Bug Resolved Urgent Scaler data missing / low count Hanjie Liu 02/19/2024 09:27 AM Actions
945 NPS software Task New High Add plots for skim files 02/15/2024 10:43 AM Actions
944 NPS software Task In Progress High Add Carbon boiling studies to BCM calibration ? 02/26/2024 10:40 AM Actions
943 NPS Task New Normal Check HV in EPICS Joshua Crafts 02/15/2024 10:44 AM Actions
942 NPS software Task In Progress Urgent Waveform analysis on DVCS data Wassim Hamdi 02/19/2024 09:27 AM Actions
941 NPS Task In Progress Normal Wiki page of on SIMC, location Joshua Crafts 02/22/2024 04:59 AM Actions
939 NPS Task New Normal HV stability plots Joshua Crafts 02/12/2024 10:03 AM Actions
938 SBS DAQ GEn RP GEMs Task New Normal Reorganize MPDs connection in VTP crates Holly Szumila-Vance 02/21/2024 10:11 AM Actions
937 SBS-DAQ-GEn-RP Task In Progress Normal Active analyzer FADC and TDC cabling William Henry 02/19/2024 07:55 AM Actions
936 SBS DAQ GEn RP GEMs Task In Progress Normal GEM polarimeter cabling Holly Szumila-Vance 02/21/2024 10:11 AM Actions
935 SBS-DAQ-GEn-RP Task In Progress Normal HCAL trigger cabling to TS Alexandre Camsonne 02/26/2024 11:04 AM Actions
934 SBS-DAQ-GEn-RP Task New Normal BB Shower VTP firmware 02/05/2024 11:14 AM Actions
933 SBS-DAQ-GEn-RP Task In Progress Normal HCAL trigger firmware Ben Raydo 02/19/2024 07:56 AM Actions
932 SBS-DAQ-GEn-RP Task In Progress Normal RF time check Alexandre Camsonne 02/28/2024 04:30 PM Actions
930 NPS Task New Normal cross check BCM vs other 02/05/2024 10:13 AM Actions
929 NPS Bug In Progress Normal FADC channel with always same signal 952 02/05/2024 09:49 AM Actions
928 NPS software Task New High Low current runs for BCM offsets Christine Ploen 02/26/2024 10:40 AM Actions
927 NPS DVCS Task In Progress High Need new calibration coefficient Hao Huang 02/08/2024 10:03 AM Actions
926 NPS Bug New Normal HMS electron trigger efficiency Mitchell Kerver 02/08/2024 10:08 AM Actions
925 NPS Bug In Progress High Scaler error Hanjie Liu 02/01/2024 09:54 AM Actions
(1-25/523) Per page: 25, 50, 100

Also available in: Atom CSV PDF